OpenStudy (anonymous):

what is the solution to this equation? 7-9b=-8b+8

OpenStudy (anonymous):

minus 7 from each side of the equation. 7-7-9b=-8b+8-7 So -9b=-8b+1

OpenStudy (anonymous):

Multiply by -1 to get: 9b=8b-1

OpenStudy (anonymous):

Sure you can work the rest out. gl

OpenStudy (whpalmer4):

Slightly different approach: Add 9b to each side of the equation to get \[7-9b+9b= -8b + 9b + 8\]\[7 = b+ 8\]\[b = 7 - 8 = -1\] As long as you do the same thing to both sides of the equation, you're good.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I got it right first so get the Nobel - hint hint ;D

OpenStudy (whpalmer4):

the box does say "Best Response" not "First Response" :-)

OpenStudy (anonymous):

Yes I agree.

OpenStudy (anonymous):

I gave BOTH ;D tehehehehe